# Self-Supervised Learning of Time Series Representation via Diffusion Process and Imputation-Interpolation-Forecasting Mask

### v1

@@ 1. Introduction
Due to their advanced generative capabilities, diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020a; Song and Ermon, 2020; Song et al., 2021a, b) have emerged as a promising solution for TS modeling, adept at handling the complexities and long-term dependencies often found in MTS data. While these methods have shown success in specific tasks like forecasting (Rasul et al., 2021a) and imputation (Tashiro et al., 2021), their adoption in SSL TSRL remains largely unexplored, leaving a gap in the related research literature. Our work, TSDE (Time Series Diffusion Embedding), pioneers in this area by integrating conditional diffusion processes with crossover Transformer encoders and introducing a Imputation-Interpolation-Forecasting (IIF) mask strategy. This unique combination allows TSDE to generate versatile representations that are applicable to a wide range of tasks, including imputation, interpolation, forecasting, classification, anomaly detection, and clustering. Our main contributions are:
@@ 2. Related work
Reconstructive methods focus on minimizing the discrepancy between original and reconstructed MTS data, mostly using an encoder-decoder Neural Network (NN) architecture to emphasize salient features and filter out noise, thereby training the NN to learn meaningful representations (Hinton and Salakhutdinov, 2006). Recent mainstream methods in this category predominantly employ CNN (Convolutional NN) (Song et al., 2020; Zhang and Liu, 2021), RNN (Recurrent NN) (Malhotra et al., 2017; Sagheer and Kotb, 2019) or Transformer (Chowdhury et al., 2022; Zhang et al., 2023b) as their architectural backbone. In this category, deep clustering stands out by simultaneously optimizing clustering and reconstruction objectives. It has been implemented through various clustering algorithms, including $k$-means(Xie et al., 2016; Caron et al., 2018), Gaussian Mixture Model (GMM) (Jiang et al., 2017; Cao et al., 2020), and spectral clustering (Tao et al., 2020). Reconstructive methods might face limitations in addressing long-term dependencies and adequately representing complex features such as seasonality, trends, and noise in extensive, high-dimensional datasets.
@@ 3.1.1. Forward process
which is a conditional Gaussian distribution with a mean of $\small\smash{\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}}$ and a covariance matrix of $\beta_{t}\textbf{I}$, and $\beta_{t}\!\in\!(0,1)$ indicates the noise level at each diffusion step $t$. Because of the properties of Gaussian kernels, we can sample any $\mathbf{x}_{t}$ from $\mathbf{x}_{0}$ directly with
@@ 7       $l^{\prime}\leftarrow$ uniformly sample a time step from $\mathbb{Z}\cap[1,L]$;
3.4. Training Loss and Procedure
The self-supervised and mini-batch training procedure, detailed in Algorithm 2 essentially attempts to solve $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\widetilde{\mathcal{L}}(\boldsymbol{\theta},\boldsymbol{\phi})$. In each iteration $i$ of the training process, a random diffusion step $t$ is chosen, at which point the denoising operation is applied.
@@ 3.5. Embedding Function
3.6. The Overall Architecture
3.7. Downstream Tasks and Model Efficiency
(a) Imputation on PhysioNet
(b) Interpolation on Electricity
(c) Forecasting on Electricity
Figure 2. Comparison of predicted and ground truth values for (a) imputation (10% missing), (b) interpolation, and (c) forecasting. The line is the median of the predictions and the red shade indicates 5$\sim$95quantile for missing/future values. See Appendix E for more results.
Datasets CSDI* (sec.) TSDE (sec.)
Electricity 1,997 163
Solar 608 62
Taxi 27,533 1,730
Traffic 7,569 422
Wiki 9,138 391
4. Experiments
@@ 4.1.1. Imputation
We carry out imputation experiments on PhysioNet [^3] PhysioNet, a healthcare dataset with 4,000 records of 35 variables over 48 hours, is processed and hourly sampled as (Salinas et al., 2019; Tashiro et al., 2021), leading to $\sim$80missing rate. For testing, we randomly mask 10%, 50%, and 90% of observed values to create ground-truth scenarios. On this dataset, we pretrain TSDE for 2,000 epochs, followed by a 200-epoch finetuning with an imputation mask. (Silva et al., 2012) and PM2.5444PM2.5, an air quality dataset, features hourly readings from 36 Beijing stations over 12 months with artificially generated missing patterns. Adapting (Tashiro et al., 2021), each series spans 36 consecutive timestamps. On this dataset, we pretrain for 1,500 epochs and finetune for 100 epochs using a history mask as detailed in Algorithm 5. (Yi et al., 2016). TSDE is benchmarked against several state-of-the-art TS imputation models. These include BRITS (Cao et al., 2018), a deterministic method using bi-directional RNN for correlation capture; V-RIN (Mulyadi et al., 2022), employing variational-recurrent networks with feature and temporal correlations for uncertainty-based imputation; GP-VAE (Fortuin et al., 2020), integrating Gaussian Processes with VAEs; and CSDI (Tashiro et al., 2021), the top-performing model among the diffusion-based TS imputation models. The model performance is evaluated using CRPS (continuous ranked probability score) to assess the fit of predicted outcomes with original data distributions, and two deterministic metrics – MAE (mean absolute error) and the RMSE (root mean square error). Deterministic metrics are calculated using the median across all samples, and CRPS value is reported as the normalized average score for all missing values distributions (approximated with 100 samples).
@@ 4.1.3. Forecasting
Our forecasting experiments employ five real-world datasets: (1) Electricity, tracking hourly consumption across 370 customers; (2) Solar, detailing photovoltaic production at 137 Alabama stations; (3) Taxi, recording half-hourly traffic from 1,214 New York locations; (4) Traffic, covering hourly occupancy rates of 963 San Francisco car lanes; and (5) Wiki, monitoring daily views of 2,000 Wikipedia pages. Adapting the practices from (Salinas et al., 2019; Nguyen and Quanz, 2021; Tashiro et al., 2021), each dataset is converted into a series of multivariate sequences, with $L_{1}$ historical timestamps followed by $L_{2}$ timestamps for forecasting. Training data apply a rolling window approach with a stride of 1, while validation and testing data employ a stride of $L_{2}$, ensuring distinct, non-overlapping series for evaluation. Specific $L_{1}$ and $L_{2}$ values are outlined in Table  8. For evaluation metrics, we use CRPS and MSE, supplemented by CRPS-Sum, as introduced in (Salinas et al., 2019). CRPS-Sum is computed by summing across different features, capturing the joint impact of feature distributions. As of benchmarking baselines, we include several state-of-the-art probabilistic MTS forecasting models: GP-copula (Salinas et al., 2019), TransMAF (Rasul et al., 2021b) and TLAE (Nguyen and Quanz, 2021). Additionally, in the realm of diffusion-based methods, we include CSDI (Tashiro et al., 2021) and TimeGrad (Rasul et al., 2021a).
@@
The forecasting results, as detailed in Table 3, showcase TSDE’s robust performance, especially when finetuned with a forecasting mask. Its effectiveness is notable when compared to CSDI, which is the most closely related method, sharing a diffusion backbone. TSDE particularly excels in the Electricity, Taxi, and Wiki datasets, especially as evaluated by the CRPS metric. However, it is important to note a discrepancy in the Solar dataset performance between TSDE/CSDI and other baselines, likely due to a data split issue: the actual test set, per the source code, is identical to the training set, which contradicts the details reported in the corresponding paper. For a qualitative illustration, refer to Figure 2(c).
@@ 4.1.5. Inference Efficiency
4.2. Anomaly Detection
In this task, TSDE is benchmarked against an extensive set of baselines featuring diverse backbones, including a) Frozen pretrained LLM-based models: GPT4TS (Zhou et al., 2023); b) Task-agnostic foundation models: TimesNet (Wu et al., 2023); c) MLP (multi-layer perceptron) based models: LightTS (Zhang et al., 2022) and DLinear (Zeng et al., 2023); and finally d) Transformer-based models: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2022b), LogSparse Transformer (Li et al., 2019), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022a), ETSformer (Woo et al., 2023), PatchTST (Nie et al., 2023) and Anomaly Transformer (Xu et al., 2022). The results in Table  6 reveal that TSDE’s anomaly detection performance surpasses nearly all baselines, with less than a 1% F1 score difference from GPT4TS. Notably, while TSDE doesn’t outperform GPT4TS, it’s important to consider that GPT4TS benefits from a pretrained LLM (GPT-2) with about 1.5 billion model parameters. TSDE, in contrast, relies on just two single-layer Transformer encoders (¡0.3 million parameters), demonstrating its competitive edge despite having significantly fewer model parameters.
@@ 4.3. Classification
As shown in Table  6, TSDE surpasses all existing baselines and is on par with the state-of-the-art BRITS baseline. It is worth noting that BRITS achieves that performance by employing a sophisticated multi-task learning mechanism tailored for classification tasks. In contrast, our method achieves top-tier results by simply finetuning a simple MLP. TSDE’s remarkable performance, especially in challenging classification scenarios with significant class imbalance ($\sim$10positive classes), highlights its ability to learn generic embeddings well-suited for downstream MTS classification tasks.
@@ 5. Conclusion
In this paper, we propose TSDE, a novel SSL framework for TSRL. TSDE, the first of its kind, effectively harnesses a diffusion process, conditioned on an innovative dual-orthogonal Transformer encoder architecture with a crossover mechanism, and employs a unique IIF mask strategy. Our comprehensive experiments across diverse TS analysis tasks, including imputation, interpolation, forecasting, anomaly detection, classification, and clustering, demonstrate TSDE’s superior performance compared to state-of-the-art models. Specifically, TSDE shows remarkable results in handling MTS data with high missing rates and various complexities, thus validating its effectiveness in capturing the intricate MTS dynamics. Moreover, TSDE not only significantly accelerates inference speed but also showcases its versatile embeddings through qualitative visualizations, encapsulating key MTS characteristics. This positions TSDE as a robust, efficient, and versatile advancement in MTS representation learning, suitable for a wide range of MTS tasks. Future work will focus on enabling the pretrained TSDE to execute IIF tasks independently, without the need for the diffusion block.
@@ A.3. Anomaly Detection datasets
Each dataset is preprocessed into a set of MTS using a non-overlapping sliding window of size 100, ensuring uniformity in the analysis. Below, we describe the five datasets employed in our experiments. The specifications of each dataset are summarized in Table  9.

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671673isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Unsupervised learningccs: Mathematics of computing Time series analysisccs: Mathematics of computing Probabilistic algorithmsccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Due to their advanced generative capabilities, diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020a; Song and Ermon, 2020; Song et al., 2021a, b) have emerged as a promising solution for TS modeling, adept at handling the complexities and long-term dependencies often found in MTS data. While these methods have shown success in specific tasks like forecasting (Rasul et al., 2021a) and imputation (Tashiro et al., 2021), their adoption in SSL TSRL remains largely unexplored, leaving a gap in the related research literature. Our work, Time Series Diffusion Embedding (TSDE), pioneers in this area by integrating conditional diffusion processes with crossover Transformer encoders and introducing an Imputation-Interpolation-Forecasting (IIF) mask strategy. This unique combination allows TSDE to generate versatile representations that are applicable to a wide range of tasks, including imputation, interpolation, forecasting, classification, anomaly detection, and clustering. Our main contributions are:
@@ 2. Related work
Reconstructive methods focus on minimizing the discrepancy between original and reconstructed MTS data, mostly using an encoder-decoder Neural Network (NN) architecture to emphasize salient features and filter out noise, thereby training the NN to learn meaningful representations (Hinton and Salakhutdinov, 2006). Recent mainstream methods in this category predominantly employ Convolutional NN (CNN) (Song et al., 2020; Zhang and Liu, 2021), Recurrent NN (RNN) (Malhotra et al., 2017; Sagheer and Kotb, 2019) or Transformer (Chowdhury et al., 2022; Zhang et al., 2023b) as their architectural backbone. In this category, deep clustering stands out by simultaneously optimizing clustering and reconstruction objectives. It has been implemented through various clustering algorithms, including $k$-means (Xie et al., 2016; Caron et al., 2018), Gaussian Mixture Model (GMM) (Jiang et al., 2017; Cao et al., 2020), and spectral clustering (Tao et al., 2020). Reconstructive methods might face limitations in addressing long-term dependencies and adequately representing complex features such as seasonality, trends, and noise in extensive, high-dimensional datasets.
@@ 3.1.1. Forward process
which is a conditional Gaussian distribution with a mean of
$\small\smash{\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}}$ and a covariance matrix of $\beta_{t}\textbf{I}$, and $\beta_{t}\!\in\!(0,1)$ indicates the noise level at each diffusion step $t$. Because of the properties of Gaussian kernels, we can sample any $\mathbf{x}_{t}$ from $\mathbf{x}_{0}$ directly with
@@ 3.3. Conditional Reverse Diffusion Process
3.4. Training Loss and Procedure
The self-supervised and mini-batch training procedure, detailed in Algorithm 2, essentially attempts to solve $\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\widetilde{\mathcal{L}}(\boldsymbol{\theta},\boldsymbol{\phi})$. In each iteration $i$ of the training process, a random diffusion step $t$ is chosen, at which point the denoising operation is applied.
3.5. Embedding Function
@@ 7       $l^{\prime}\leftarrow$ uniformly sample a time step from $\mathbb{Z}\cap[1,L]$;
3.6. The Overall Architecture
3.7. Downstream Tasks and Model Efficiency
@@ 4.1.1. Imputation
We carry out imputation experiments on PhysioNet [^3] PhysioNet, a healthcare dataset with 4,000 records of 35 variables over 48 hours, is processed and hourly sampled as (Salinas et al., 2019; Tashiro et al., 2021), leading to $\sim$80missing rate. For testing, we randomly mask 10%, 50%, and 90% of observed values to create ground-truth scenarios. On this dataset, we pretrain TSDE for 2,000 epochs, followed by a 200-epoch finetuning with an imputation mask. (Silva et al., 2012) and PM2.5444PM2.5, an air quality dataset, features hourly readings from 36 Beijing stations over 12 months with artificially generated missing patterns. Adapting (Tashiro et al., 2021), each series spans 36 consecutive timestamps. On this dataset, we pretrain for 1,500 epochs and finetune for 100 epochs using a history mask as detailed in Algorithm 5. (Yi et al., 2016). TSDE is benchmarked against several state-of-the-art TS imputation models. These include BRITS (Cao et al., 2018), a deterministic method using bi-directional RNN for correlation capture; V-RIN (Mulyadi et al., 2022), employing variational-recurrent networks with feature and temporal correlations for uncertainty-based imputation; GP-VAE (Fortuin et al., 2020), integrating Gaussian Processes with VAEs; and CSDI (Tashiro et al., 2021), the top-performing model among the diffusion-based TS imputation models. The model performance is evaluated using continuous ranked probability score (CRPS) to assess the fit of predicted outcomes with original data distributions, and two deterministic metrics – mean absolute error (MAE) and the root mean square error (RMSE). Deterministic metrics are calculated using the median across all samples, and CRPS value is reported as the normalized average score for all missing values distributions (approximated with 100 samples).
@@ 4.1.3. Forecasting
We conducted two sets of benchmarking experiments. The first was a benchmarking for probabilistic multivariate time series forecasting. We employ five real-world datasets: (1) Electricity, tracking hourly consumption across 370 customers; (2) Solar, detailing photovoltaic production at 137 Alabama stations; (3) Taxi, recording half-hourly traffic from 1,214 New York locations; (4) Traffic, covering hourly occupancy rates of 963 San Francisco car lanes; and (5) Wiki, monitoring daily views of 2,000 Wikipedia pages. Adapting the practices from (Salinas et al., 2019; Nguyen and Quanz, 2021; Tashiro et al., 2021), each dataset is converted into a series of multivariate sequences, with $L_{1}$ historical timestamps followed by $L_{2}$ timestamps for forecasting. Training data apply a rolling window approach with a stride of 1, while validation and testing data employ a stride of $L_{2}$, ensuring distinct, non-overlapping series for evaluation. Specific $L_{1}$ and $L_{2}$ values are outlined in Table  9. For evaluation metrics, we use CRPS and MSE, supplemented by CRPS-Sum, as introduced in (Salinas et al., 2019). CRPS-Sum is computed by summing across different features, capturing the joint impact of feature distributions. As of benchmarking baselines, we include several state-of-the-art probabilistic MTS forecasting models: GP-copula (Salinas et al., 2019), TransMAF (Rasul et al., 2021b) and TLAE (Nguyen and Quanz, 2021). Additionally, in the realm of diffusion-based methods, we include CSDI (Tashiro et al., 2021) and TimeGrad (Rasul et al., 2021a).
For the second benchmarking, which is a deterministic benchmarking including recent baselines in the time series library (Wu et al., 2023), we conducted five experiments for each baseline following the same setting in (Wu et al., 2023) with history-prediction window lengths of {8-8, 16-16, 32-32, 96-96, 96-192} on the Electricity dataset. We report the averaged performance in terms of MAE and MSE. We compared TSDE with the following baselines: TimesNet (Wu et al., 2023), ETSformer (Woo et al., 2023), LightTS (Zhang et al., 2022), DLinear (Zeng et al., 2023), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022a), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2022b), Informer (Zhou et al., 2021), Reformer (Kitaev et al., 2020), and PatchTST (Nie et al., 2023).
Table 2. Forecasting task results on Electricity following (Wu et al., 2023) setting. We compare extensive competitive models under five different history-prediction lengths {8-8, 16-16, 32-32, 96-96, 96-192}. Avg is averaged from all five history-prediction lengths results. See Table 13 for full results.
Models
TSDE
TimesNet
ETSformer
LightTS
DLinear
FEDformer
Stationary
Autoformer
Pyraformer
Informer
Reformer
PatchTST
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Avg
0.169
0.253
0.195
0.290
0.241
0.355
0.227
0.329
0.359
0.408
0.193
0.310
0.191
0.285
0.192
0.311
0.299
0.367
0.295
0.390
0.287
0.382
0.342
0.362
  • $\ast$ means that there are some mismatches between our input-output setting and their papers. We adopt their official codes and only change the length of input and output sequences for a fair comparison.
@@
The forecasting results, as detailed in Table 3, showcase TSDE’s robust performance, especially when finetuned with a forecasting mask. Its effectiveness is notable when compared to CSDI, which is the most closely related method, sharing a diffusion backbone. TSDE particularly excels in the Electricity, Taxi, and Wiki datasets, especially as evaluated by the CRPS metric. However, it is important to note a discrepancy in the Solar dataset performance between TSDE/CSDI and other baselines, likely due to a data split issue: the actual test set, per the source code, is identical to the training set, which contradicts the details reported in the corresponding paper. Table 2 demonstrates that TSDE outperforms the recent baselines in terms of average MSE and MAE, highlighting its robustness and superiority compared to recent methods. The detailed results for each window length are available in the appendix, Table 13. For a qualitative illustration, refer to Figure 2(c).
@@ 4.1.5. Inference Efficiency
(a) Imputation on PhysioNet
(b) Interpolation on Electricity
(c) Forecasting on Electricity
Figure 2. Comparison of predicted and ground truth values for (a) imputation (10% missing), (b) interpolation, and (c) forecasting. The line is the median of the predictions and the red shade indicates 5$\sim$95quantile for missing/future values. See Appendix G for more results.
Datasets CSDI* (sec.) TSDE (sec.)
Electricity 1,997 163
Solar 608 62
Taxi 27,533 1,730
Traffic 7,569 422
Wiki 9,138 391
@@
4.2. Anomaly Detection
In this task, TSDE is benchmarked against an extensive set of baselines featuring diverse backbones, including a) Frozen pretrained LLM-based models: GPT4TS (Zhou et al., 2023); b) Task-agnostic foundation models: TimesNet (Wu et al., 2023); c) MLP (multi-layer perceptron) based models: LightTS (Zhang et al., 2022) and DLinear (Zeng et al., 2023); and finally d) Transformer-based models: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2022b), LogSparse Transformer (Li et al., 2019), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022a), ETSformer (Woo et al., 2023), PatchTST (Nie et al., 2023) and Anomaly Transformer (Xu et al., 2022). The results in Table  7 reveal that TSDE’s anomaly detection performance surpasses nearly all baselines, with less than a 1% F1 score difference from GPT4TS. Notably, while TSDE doesn’t outperform GPT4TS, it’s important to consider that GPT4TS benefits from a pretrained LLM (GPT-2) with about 1.5 billion model parameters. TSDE, in contrast, relies on just two single-layer Transformer encoders (¡0.3 million parameters), demonstrating its competitive edge despite having significantly fewer model parameters.
@@ 4.3. Classification
As shown in Table  7, TSDE surpasses all existing baselines and is on par with the state-of-the-art BRITS baseline. It is worth noting that BRITS achieves that performance by employing a sophisticated multi-task learning mechanism tailored for classification tasks. In contrast, our method achieves top-tier results by simply finetuning a simple MLP. TSDE’s remarkable performance, especially in challenging classification scenarios with significant class imbalance ($\sim$10positive classes), highlights its ability to learn generic embeddings well-suited for downstream MTS classification tasks.
@@ 5. Conclusion
In this paper, we propose TSDE, a novel SSL framework for TSRL. TSDE, the first of its kind, effectively harnesses a diffusion process, conditioned on an innovative dual-orthogonal Transformer encoder architecture with a crossover mechanism, and employs a unique IIF mask strategy. Our comprehensive experiments across diverse TS analysis tasks, including imputation, interpolation, forecasting, anomaly detection, classification, and clustering, demonstrate TSDE’s superior performance compared to state-of-the-art models. Specifically, TSDE shows remarkable results in handling MTS data with high missing rates and various complexities, thus validating its effectiveness in capturing the intricate MTS dynamics. Moreover, TSDE not only significantly accelerates inference speed but also showcases its versatile embeddings through qualitative visualizations, encapsulating key MTS characteristics. This positions TSDE as a robust, efficient, and versatile advancement in MTS representation learning, suitable for a wide range of MTS tasks. Future work will focus on several key directions to address the limitation of slower inference for IIF tasks. Particularly, we will explore simplifying TSDE’s architecture with a simple MLP without the need for the diffusion block, enabling the pretrained TSDE to execute IIF tasks independently.
Acknowledgements.
We would like to thank Yang Song (OpenAI & Stanford University) for his help to connect us with CSDI authors and the insightful discussions regarding time series representation learning. R. Tu would like to acknowledge the support of Gustav Eje Henter, Hedvig Kjellström and the Wallenberg AI, Autonomous Systems and Software Program (WASP). R. Tu was also funded by the Industrial Strategic Technology Development Program (grant no. 20023495) from MOTIE, Korea.
@@ A.3. Anomaly Detection datasets
Each dataset is preprocessed into a set of MTS using a non-overlapping sliding window of size 100, ensuring uniformity in the analysis. Below, we describe the five datasets employed in our experiments. The specifications of each dataset are summarized in Table  10.
@@ D.3.1. Probabilistic Multivariate Forecasting
D.3.2. Deterministic multivariate forecasting
For the determinitic multivariate forecasting benchmarking, five experiments were conducted on the preprocessed Electricity dataset provided by TimesNet authors (Wu et al., 2023). We used the same configuration and data splits as in (Wu et al., 2023). To create the MTS, we utilized a window of length $L=L_{1}+L_{2}$, where $L_{1}$ is the history length and $l_{2}$ is the prediction length. The values employed for these lengths in our experiments for this benchmarking are: 8-8, 16-16, 32-32, 96-96, and 96-192 where the first value is $L_{1}$ and the second value is $L_{2}$. Each experiment was run only once.
The hyperparameters, such as batch size and the number of training epochs, are presented in Table 12. During training, we employed the IIF masking strategy. Larger prediction windows as in (Wu et al., 2023), could not be used with TSDE due to the out of memory limitation.
For all the other baselines from the time series library (Wu et al., 2023), we utilized the original implementations published by the authors, using the default hyperparameters provided in the experiments scripts to maintain consistenty with their original long-term forecasting experiments.
Table 12. Hyperparameters used in the experiments for different window lengths
Hyperparameters 8-8 16-16 32-32 96-96 96-192
Batch size 8 8 8 2 1
Number of training epochs 100 100 100 20 20